When 4.84 g4.84 g of a nonelectrolyte solute is dissolved in water to make 475 mL475 mL of solution at 26 °C,26 °C, the solution exerts an osmotic pressure of 867 torr.867 torr.

What is the molar concentration of the solution?

concentration:

MM

How many moles of solute are in the solution?

moles of solute:

molmol

What is the molar mass of the solute?

molar mass:

g/mol
Answer #1

1)

P= 867.0 torr

= (867.0/760) atm

= 1.1408 atm

T= 26.0 oC

= (26.0+273) K

= 299 K

use:

P = C*R*T

1.141 = C*0.08206*299

C =4.649*10^-2 M

Answer: 4.65*10^-2 M

2)

volume , V = 475 mL

= 0.475 L

use:

number of mol,

n = Molarity * Volume

= 4.649*10^-2*0.475

= 2.208*10^-2 mol

Answer: 2.21*10^-2 mol

3)

mass(solute)= 4.84 g

use:

number of mol = mass / molar mass

2.208*10^-2 mol = (4.84 g)/molar mass

molar mass = 2.192*10^2 g/mol

Answer: 219 g/mol